# HAMMER: Heterogeneous, Multi-Robot Semantic Gaussian Splatting

### v1

@@ VI Conclusion and Limitations
HAMMER is the first online multi-robot Gaussian Splatting mapping pipeline, built for heterogeneous robot teams. Maps produced with HAMMER can be used for a variety of downstream tasks such as language-guided navigation. We show that HAMMER outperforms comparable baselines on real datasets and approaches the visual fidelity of an oracle. Future work will focus on better methods for data curation, reducing the volume of redundant data, and enabling detailed scene reconstructions over larger areas. We hope to extend HAMMER to track and map dynamic objects, and leverage the semantics embedded in the HAMMER map to extract intelligent spatial relationships between objects.
HAMMER has several limitations. Devices which do not provide pose estimates cannot be used in HAMMER, and very noisy estimates cannot be rectified by HAMMER’s pose refinement and produce a poor quality map. Moreover, the current implementation of HAMMER does not account for RGB-only devices (no geometric data). However, we believe HAMMER can be extended for use in RGB-only $3$DGS SLAM, an active research area.

### v2

@@ V Conclusion and Limitations
HAMMER is the first online multi-robot 3DGS mapping pipeline, built for heterogeneous robot teams. Maps produced with HAMMER can be used for a variety of downstream tasks like language-guided navigation. We show that HAMMER outperforms comparable baselines on real datasets and approaches the visual fidelity of an oracle in less than $10\%$ of the time. Future work will focus on better methods for data curation, volume reduction of redundant data, and detailed scene reconstructions over larger areas. The flexibility of HAMMER also facilitates map constructions from a wide range of robot embodiments (e.g. Aria glasses) and investigating the benefits of fusing diverse sets of devices is an interesting avenue for future research.
Figure 6: Motion plans for robots navigating to language-specified goals with trajectories from Splat-Nav [7].
Acknowledgment
The authors would like to thank Dr. Ola Shorinwa for his valuable comments and thoughtful advice on this work.